OpenStudy (anonymous):

the "lower portion" of the greenhouse has a volume of 528 cubic yards. it has a height of x yards, a width of 5x-4 yards, and a length of 20x-44 yards. Write an equation equal to zero that will allow you to solve for x.

Directrix (directrix):

Assuming that the lower portion of the greenhouse is a rectangular solid, its volume is found by multiplying the three dimensions of the solid. V = l*w*h 528 = x (5x-4) (20x-44) 528 - x (5x-4) (20x-44) = 0